40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 25 15:30:29 2013, yuta, Update, Alignment, Y arm locked, both colors 
    Reply  Mon Feb 25 17:58:28 2013, rana, Update, Alignment, Y arm locked, both colors 
       Reply  Mon Feb 25 22:42:32 2013, yuta, Summary, Alignment, current IFO situation IPANGdrift.png
    Reply  Mon Feb 25 21:25:14 2013, yuta, Update, Alignment, both arms locked in IR QUAD1_1045890588.png
       Reply  Tue Feb 26 01:49:27 2013, Jenne, Update, Locking, PRMI locked 
          Reply  Tue Feb 26 02:04:51 2013, Koji, Update, Locking, PRMI locked 
             Reply  Tue Feb 26 11:55:11 2013, Manasa, Update, Locking, FPMI locked 
                Reply  Tue Feb 26 15:32:47 2013, Jenne, Update, Locking, FPMI locked 
             Reply  Wed Feb 27 01:09:53 2013, yuta, Update, Locking, PRMI sideband/carrier locked 6x
                Reply  Wed Feb 27 02:44:39 2013, Jenne, Update, Locking, DRFPMI flashes 
                   Reply  Wed Feb 27 20:10:43 2013, rana, Update, Locking, DRFPMI flashes 
Message ID: 8162     Entry time: Mon Feb 25 21:25:14 2013     In reply to: 8157     Reply to this: 8165
Author: yuta 
Type: Update 
Category: Alignment 
Subject: both arms locked in IR 

[Jenne, Evan, Yuta]

After Y alignment, X arm is aligned to IR and we got both arms locked in IR.
There's some dift (input pointing?) and this made aligning both arms tough. I will elog about it later.
Attached is ETMYF. ETMXF, ITMYF, ITMXF when both arms are locked by IR.

Alignment Procedure:
  1. Algined BS/ITMX to get MI fringe in AS. We got X arm flashing at this point.
  2. Use BS/ITMX/ETMX to get TRX maximized, without losing good MI fringe in AS. We reached 0.75.
  3. There was clipping of TRX beam at Xend optics. Since whole IFO alignment is started from Y green, this clipping is because of poor Y green pointing. But we needed clear TRX for aligning Xarm, so we re-arranged Xend TRX path to avoid clipping.

X arm issues:
  - Beam motion at TRX is larger than TRY. Turning off clean table air didn't help. Maybe we need BS oplev on or ETMX coil gain balancing.
  - X green scatters into TRX PD and ETMXT camera. Fix them.

Attachment 1: QUAD1_1045890588.png  325 kB  | Hide | Hide all
QUAD1_1045890588.png
ELOG V3.1.3-